OpenStudy (anonymous):

Which logarithmic graph can be used to approximate the value of y in the equation 2^y = 3?

OpenStudy (anonymous):

@johnweldon1993 @kohai @KyanTheDoodle @Loser66

OpenStudy (anonymous):

@sasogeek

OpenStudy (anonymous):

@Luigi0210 @Whitemonsterbunny17

OpenStudy (kyanthedoodle):

Honestly, I have no idea.

OpenStudy (anonymous):

it's alright @KyanTheDoodle , thank you

OpenStudy (anonymous):

If the y was an x, do you know what the graph would look like?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Let's identify what we know. Is 2^y ever negative? Is it ever 0?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

What happens as y gets larger? What happens when y is negative?

OpenStudy (anonymous):

if it's negative the graph goes down i think and the exponent determines which way it goes,

OpenStudy (anonymous):

With a larger and larger negative exponent, it gets closer and closer to 0. With a larger and larger positive exponent, it gets larger and larger. So, we're looking for the graph that never results in a negative x value. It gets closer to 0 with the larger negative y, and gets further and further from the y axis as y gets larger. Which of the graphs follows this pattern?

OpenStudy (anonymous):

I eliminated B and c because they look they're getting close to zero and D is increasing in y value but to the negative so i think the answer is "A"

OpenStudy (anonymous):

A goes into the negative x values, so it can't be A. So does D.

OpenStudy (anonymous):

Is B increasing in y value the most?

OpenStudy (anonymous):

I'm not sure what you mean by that. We're looking for the graph of the equation: 2^y We know what this will do in certain situations (such as it will never go negative). This means we can eliminate graphs A and D, as the line shown on there goes into negative X values.

OpenStudy (anonymous):

After that, we can look at 2^1 = 2 and 2^0 = 1 to eliminate one of the other two graphs.

OpenStudy (anonymous):

why did you have them to the ^1 and ^0

OpenStudy (anonymous):

Those are the results when y = 1 and y = 0. We can see which graph of the remaining two goes through the correct points. y = 0: 2^0 = 1 So we look for the point where x=1 and y=0, and see if the graph goes through that point.

OpenStudy (anonymous):

i see that in C

OpenStudy (anonymous):

i'm sorry for taking long my internet connection was slow

OpenStudy (anonymous):

@Vandreigan

OpenStudy (anonymous):

Yep, C. It follows all the patterns :)

OpenStudy (anonymous):

alright, thank you so much for your help.. i appreciate it @Vandreigan

OpenStudy (anonymous):

My pleasure :)